39. The Puppet Co. has the following unit and mix data: Dah To $4.00 0.6 20% $5.00 Unit sales price Unit contribution margin Sales mix (S) Fixed costs Target 0.75 $99,000 24.750 How many units of Dah must be sold at the breakeven point? a) 75,000 b) 27,500 c) 37,500 d) 55,000 Answer: b Difficulty: Easy Learning Objective: Apply CVP calculations for a single CPA: Management Accounting 40. The Puppet Co. has the following unit and mix data: Dah To $4.00 20% $99,000 24,750 $5.00 Unit sales price Unit contribution margin Sales mix ($) Fixed costs Target profit 0.75 80% 0.60 How many units in total must be sold to earn the target profit? a) 687,500 b) 165,000 c) 112,500 d) 171,875 Answer: d Difficulty: Easy Learning Objective: Apply CVP calculations for a single
41. MacLean Company produces a single product. Following is cost information: Variable Cost Fixed Costs per Unit Manufacturing costs Non-manufacturing costs $35,000 $15 60,000 If the product sells for $60, how many units must be sold to break even? a) 1,000 b) 2,375 c) 2,714 d) 3,800 Answer: Difficulty: Easy Learning Objective: Apply CVP calculations for a single product. CPA: Management Accounting 42. Ciao Company manufactures a single product. The product sells for $10. The variable manufacturing cost per unit is $2 and the variable selling cost is $2 per unit. Ciao incurs monthly fixed costs of $100,000 for manufacturing and 140,000 for administration and selling. Ciao Company is considering changes to its production and distribution procedures. If the changes are made, total variable costs (manufacturing and selling) will be $3 and total fixed costs (manufacturing, administration, and selling) will be $350,000 per month. The selling price will remain at $10.Н the changes are made, the number of units required to break even will be: a) Greater than before b) The same as before c) Less than before d) Cannot be determined Answer: a Difficulty: Easy Learning Objective: Apply CVP calculations for a single product. CPA: Management Accounting 43. The breakeven point can be defined as:
Answer #1

Please provide rating..

Answer 39
Do Dah total
unit sale price 5 4
Unit contribution margin 0.75 0.6
Mix 80% 20%
Unit contribution margin * mix          0.60          0.12               0.72
Fixed cost = 99000
Break even unit = 99000/.72        137,500
Dah must sale = 137500*20%           27,500
Correct answer is option b. 27500
Answer 40
Do Dah total
unit sale price 5 4
Unit contribution margin 0.75 0.6
Mix 80% 20%
Unit contribution margin * mix          0.60          0.12               0.72
Fixed cost = 99000
Target profit = 24750
Total contribution required 123750
Unit required to sale = 123750/.72        171,875
Correct answer is option d. 171875
Answer 41
Sale price per unit 60
Variable cost per unit =15+10 25
Contribution margin 35
Fixed cost =35000+60000 95000
Break even unit =95000/35        2,714
Correct answer is option c. 2714
Answer 42
Current break even unit
Sale price per unit 10
Variable cost per unit =2+2 4
Contribution margin 6
Fixed cost =100000+140000 240000
Break even unit =240000/6      40,000
New break even unit
Sale price per unit 10
Variable cost per unit =2+2 3
Contribution margin 7
Fixed cost 350000
Break even unit =240000/6      50,000
Correct answer is option a. Greater than before (50000 unit compared to 40000 earlier)